# Parametric Phase Modulation in Superconducting Circuits

### v1

@@ Abstract
Parametric modulation is widely employed in superconducting circuits for quantum simulations and high-fidelity two-qubit gates, valued for its versatility. Conventionally, the qubit coupling strength is determined by the amplitude of the parametric flux pulse, which affects qubit parameters dramatically. In this article, we propose and implement a phase modulation scheme to tune the interaction strength via adjusting the relative phase between the parametric flux pulses applied to two coupled qubits. We characterize this modulation for sideband couplings, at both sweet and off-sweet spots, achieving a broad range of coupling strengths as confirmed by both population dynamics and spectroscopy methods. This approach enables phase-controlled modulation of coupling strength, providing a promising candidate for parametrically driven quantum simulations and gate operations.
@@ I Introduction
The superconducting quantum circuits have emerged as a leading platform for large-scale quantum simulation and computation due to their high controllability, flexibility, and scalability [1]. To enable the scaling of superconducting qubits, various tunable coupling schemes have been proposed to address the challenges of large-scale systems, where couplings between qubits must be precisely tuned to suppress unwanted interactions or enhance desired ones [2, 3, 4].
@@
To meet this demand, various schemes for tunable coupling have been developed. Tunable couplers, often realized as additional circuit elements, are widely implemented in transmon-based architectures to mediate qubit-qubit interactions, allowing couplings to be turned on or off and mitigating issues like parasitic coupling, frequency crowding, control crosstalk, and leakage to non-computational states [4, 5]. These advantages have facilitated the realization of large-scale quantum simulations and high-fidelity two-qubit gates [6, 7].
@@
Complementary to these hardware-based couplers, some tunable coupling schemes exploit inherent properties of superconducting qubits, such as parametric modulation [8, 9, 10] and all-microwave schemes [11, 12]. Two-qubit gates activated by parametric modulation have gained increasing attention due to their robustness against flux distortions, noise, and crosstalk [13, 10, 14, 15, 16, 17, 18, 19, 20, 21]. Periodic modulation provides net-zero and refocusing effects [22, 23], which lead to dynamical sweet spots [24, 25, 26, 27, 28, 29, 30], the continuous version of the dynamical decoupling scheme [23], and enhanced qubit coherence [31, 32, 33] (see Appendix G for details). Parametric pulses can bridge energy gaps between the far-detuned qubit states, inducing sideband interactions that help avoid frequency collisions, thus offering great flexibility for quantum simulation and computation [15, 34, 19, 35, 36, 37, 38]. This modulation enables transitions between off-resonance qubits with tunable coupling, even for parametric-resonance qubits [19].
@@
In conventional parametric modulation schemes, the coupling strength is governed by the amplitude and frequency of the applied pulses. However, tuning these parameters typically induces substantial shifts in the qubit frequencies, thereby imposing significant challenges on the calibration and scalability of large-scale quantum processor architectures. To overcome this limitation, we introduce a approach to control the interaction strength via adjusting the phase of the parametric modulation. Normally, the parametric phase is utilized in specific quantum simulation protocols for tailoring Hamiltonians [39, 40, 35, 41, 42, 43] or for dynamic on/off switching of couplings under particular resonance conditions [44, 45, 46]. Achieving precise phase control and compensation remains a significant challenge in both quantum gate operations [15, 19, 37] and simulations [35, 38] based on parametric modulation.
@@
In this work, we re-examine the conventional role of parametric phase, treating it not as an experimental overhead requiring compensation, but as a valuable and versatile quantum control resource. We propose and experimentally demonstrate a new approach where the relative phase between two simultaneous parametric flux pulses—one applied to each of two interacting superconducting transmon qubits—provides a direct control parameter for their coupling strength. A crucial advantage of this phase-based method is that it allows for precise modulation of the interaction strength without inducing the time-averaged qubit frequency shifts that typically necessitate cumbersome recalibration.
@@
We experimentally validate this mechanism in a versatile system of two qubits coupled via a tunable coupler. This configuration allows us to retain the inherent advantages of couplers while simultaneously enhancing the generality and flexibility of our parametric phase modulation scheme. Our experiments, employing population dynamics and spectroscopy methods, systematically characterize this phase-controlled interaction at both sweet and off-sweet spots, thereby demonstrating its broad applicability under diverse operating conditions.
@@ II.1 Parametric modulation
In practical frequency-tunable transmon qubits, the relationship between the qubit frequency and externally applied flux biases of qubits is nonlinear. An Arbitrary Waveform Generator (AWG) generates a programmed voltage pulse $V(t)=\bar{V}+\tilde{V}\cos(\omega_{p}t+\phi_{p})$, which is applied to the qubit, and the modulation of the flux bias pulse is described as:
@@
where the flux oscillates around the parking flux bias $\bar{\Phi}$ with parametric amplitude $\tilde{\Phi}$, frequency $\omega_{p}$, and phase $\phi_{p}$. The qubit frequency becomes time-dependent and can be expressed as the Fourier series:
@@
where the Fourier coefficients are given by $f_{k}=\frac{\omega_{p}}{\pi(1+\delta_{k,0})}\int_{0}^{\frac{2\pi}{\omega_{p}}}dt\cos[k(\omega_{p}t+\phi_{p})]\omega(t)$. The frequency can be approximated as:
@@
when the parking flux bias $\bar{\Phi}$ is set at the sweet spot of qubit and $\tilde{\Phi}\lesssim\Phi_{0}/2$ (with $\Phi_{0}=h/2e$ being the magnetic flux quantum). Here, $\bar{\omega}$ denotes the time-averaged qubit frequency under modulation, while $\epsilon_{p}$ represents the qubit frequency excursion (i.e., the amplitude of the frequency modulation). If the parking flux bias $\bar{\Phi}$ is set at the off-sweet spot, the frequency can be approximated as:
@@
where $\omega_{1,2},~\alpha_{1,2},~b_{1,2}$, $b^{\dagger}_{1,2}$, and $g$ represent the frequencies, anharmonicities, annihilation and creation operators, and bare coupling strength of qubits, respectively. When the parking flux bias $\bar{\Phi}_{1}$ is set at the off-sweet spot with a weak parametric amplitude $\tilde{\Phi}_{1}$, modulating $Q_{1}$ as $\omega_{1}(t)\approx\bar{\omega}_{1}+\epsilon_{p1}\cos(\omega_{p1}t+\phi_{p1})$, it induces frequency modulation sidebands, and the qubit frequency oscillates at many harmonics of the parametric frequency $\omega_{p1}$ (see Appendix B and E for details). In principle, all parameters $\omega_{1,2},~\alpha_{1,2}$, and $g$ under modulation in the frequency domain become time-dependent due to the interaction [9, 19, 47]. To obtain the effective Hamiltonian in a rotating frame, we define the unitary rotation transformation as $U=e^{-i\int_{0}^{t}\mathcal{H}_{0}(\tau)d\tau}.$ For simplicity, we also define $F_{1}(t)=\int_{0}^{t}\omega_{1}(\tau)d\tau$$F_{2}(t)=\int_{0}^{t}\omega_{2}(\tau)d\tau$, $A_{1}(t)=\int_{0}^{t}\alpha_{1}(\tau)d\tau$, and $A_{2}(t)=\int_{0}^{t}\alpha_{2}(\tau)d\tau$. The effective Hamiltonian is then given by
$\begin{split}H_{\mathrm{eff}}&=i\frac{dU^{\dagger}}{dt}U+U^{\dagger}(\mathcal{H}_{0}+\mathcal{H}_{I})U\\ &=g[b_{1}b_{2}e^{i(-F_{1}-A_{1}(b_{1}^{\dagger}b_{1}-I)-F_{2}-A_{2}(b_{2}^{\dagger}b_{2}-I))}\\ &+b_{1}^{\dagger}b_{2}e^{i(F_{1}+A_{1}b_{1}^{\dagger}b_{1}-F_{2}-A_{2}(b_{2}^{\dagger}b_{2}-I))}\\ &+b_{1}b_{2}^{\dagger}e^{i(-F_{1}-A_{1}(b_{1}^{\dagger}b_{1}-I)+F_{2}+A_{2}b_{2}^{\dagger}b_{2})}\\ &+b_{1}^{\dagger}b_{2}^{\dagger}e^{i(F_{1}+A_{1}b_{1}^{\dagger}b_{1}+F_{2}+A_{2}b_{2}^{\dagger}b_{2})}],\end{split}$ (6)
where we approximate $g$ as constant. In the $\{|01\rangle,|10\rangle\}$ subspace (indexing the states of coupled qubits $|Q_{1}Q_{2}\rangle$), the above effective Hamiltonian Eq. (6) is truncated as
$\begin{split}H_{\mathrm{eff}}&=ge^{i\Delta t}e^{i\frac{\epsilon_{p1}}{\omega_{p1}}[\sin(\phi_{p1})-\sin(\omega_{p1}t+\phi_{p1})]}|10\rangle\langle 01|+\text{H.c.}\\ &=g\sum_{n=-\infty}^{\infty}J_{n}(\frac{\epsilon_{p1}}{\omega_{p1}})e^{i[(\Delta+n\omega_{p1})t+\beta_{n}]}|10\rangle\langle 01|+\text{H.c.},\end{split}$ (7)
where we use the Jacobi-Anger expansion $e^{iz\sin\theta}=\sum_{n=-\infty}^{\infty}J_{n}(z)e^{in\theta}$, with $J_{n}$ being the $n$-th Bessel function of the first kind, and $\Delta=\bar{\omega}_{2}-\bar{\omega}_{1},\beta_{n}=n(\phi_{p1}+\pi)+\frac{\epsilon_{p1}}{\omega_{p1}}\sin(\phi_{p1})$. The induced effective coupling strength is given by
$g_{\text{eff}}^{n}=gJ_{n}(\frac{\epsilon_{p1}}{\omega_{p1}}),$ (8)
when the resonance condition $\Delta+n\omega_{p1}=0,n\in\mathbb{Z}$ is satisfied.
@@ II.2 Parametric phase modulation
Here, we observe that the parametric phase $\phi_{p1}$ does not affect the effective coupling strength $g_{\mathrm{eff}}^{n}$ in Eq. (8). However, introducing an additional modulation at $Q_{2}$ with a weak amplitude $\tilde{\Phi}_{2}$, given by $\omega_{2}(t)\approx\bar{\omega}_{2}+\varepsilon_{p2}\cos(\omega_{p2}t+\phi_{p2})$, renders the relative phase $\delta\phi_{p}=\phi_{p1}-\phi_{p2}$ as a modulating factor for the effective coupling strength. According to Eq. (6) and (7), the new phase-tunable coupling strength is
@@
We use sum-to-product identities followed by the Jacobi-Anger expansion to simplify the above expression in Eq. (9), rather than applying the separate Jacobi-Anger expansion [39, 40, 35, 46, 42, 43]. The coupling strength in Eq. (9) becomes time-independent
@@
when $\omega_{p}=\omega_{p1}=\omega_{p2}$ and $\Delta+n\omega_{p}=0,n\in\mathbb{Z}$, which results in $A=\mathrm{sgn}[-\sin\delta\phi_{p}/2]\sqrt{(\frac{\epsilon_{p1}}{\omega_{p}}-\frac{\epsilon_{p2}}{\omega_{p}}\cos\delta\phi_{p})^{2}+(\frac{\epsilon_{p2}}{\omega_{p}}\sin\delta\phi_{p})^{2}}$.
Additionally, the coupling strength $g_{\text{phase}}^{n}$ depends not only on parametric amplitudes $\epsilon_{p1},\epsilon_{p2}$ and frequencies $\omega_{p1},\omega_{p2}$, but also on phases $\phi_{p1},\phi_{p2}$. This indicates that the parametric relative phase $\delta\phi_{p}$ can be modulated to adjust the argument $A$ of the $n$-th Bessel function, thereby modulating the coupling strength. The range of $|A|\in\left[\left|\frac{\epsilon_{p1}}{\omega_{p1}}-\frac{\epsilon_{p2}}{\omega_{p2}}\right|,\frac{\epsilon_{p1}}{\omega_{p1}}+\frac{\epsilon_{p2}}{\omega_{p2}}\right]$ determines the strength range, which is analogous to interference effects. When the parking flux bias $\bar{\Phi}$ is set at the sweet spot with the parametric amplitude $\tilde{\Phi}\lesssim\Phi_{0}/2$, the form of the effective coupling remains the same as in Eq. (9) but with the parametric frequency $\omega_{p1}(\omega_{p2})$ and phase $\phi_{p1}(\phi_{p2})$ doubled.

### v2

@@ Abstract
Parametric modulation, valued for its versatility, is widely employed in superconducting circuits for quantum simulations and high-fidelity two-qubit gates. Conventionally, the qubit coupling strength is determined by the amplitude of the parametric flux pulse, which affects qubit parameters dramatically. In this article, we propose and implement a phase-modulation scheme to tune the interaction strength via adjustment of the relative phase between the parametric flux pulses applied to two coupled qubits. We characterize this modulation for sideband couplings, at both sweet and off-sweet spots, achieving a broad range of coupling strengths, as confirmed by both population dynamics and spectroscopy methods. This approach enables phase-controlled modulation of coupling strength, providing a promising candidate for parametrically driven quantum simulations and gate operations.
@@ I Introduction
Superconducting quantum circuits have emerged as a leading platform for large-scale quantum simulation and computation due to their high controllability, flexibility, and scalability [1]. To enable the scaling of superconducting qubits, various tunable coupling schemes have been proposed to address the challenges of large-scale systems, where couplings between qubits must be precisely tuned to suppress unwanted interactions or enhance desired ones [2, 3, 4].
@@
To meet this demand, various schemes for tunable coupling have been developed. Tunable couplers, often realized as additional circuit elements, are widely implemented in transmon-based architectures to mediate qubit-qubit interactions, allowing couplings to be turned on or off, and mitigating issues like parasitic coupling, frequency crowding, control crosstalk, and leakage to non-computational states [4, 5]. These advantages have facilitated the realization of large-scale quantum simulations and high-fidelity two-qubit gates [6, 7].
@@
Complementary to these hardware-based couplers, some tunable coupling schemes exploit the inherent properties of superconducting qubits, such as parametric modulation [8, 9, 10] and all-microwave schemes [11, 12]. Two-qubit gates activated by parametric modulation have gained increasing attention due to their robustness against flux distortions, noise, and crosstalk [13, 10, 14, 15, 16, 17, 18, 19, 20, 21]. Periodic modulation provides net-zero and refocusing effects [22, 23], which lead to dynamical sweet spots [24, 25, 26, 27, 28, 29, 30], the continuous version of the dynamical decoupling scheme [23], and enhanced qubit coherence [31, 32, 33] (see Appendix A for details). Parametric pulses can bridge the energy gaps between the far-detuned qubit states, inducing sideband interactions that help avoid frequency collisions, thus offering great flexibility for quantum simulation and computation [15, 34, 19, 35, 36, 37, 38]. This modulation enables transitions between off-resonance qubits with tunable coupling, even for parametric-resonance qubits [19].
@@
In conventional parametric modulation schemes, the coupling strength is governed by the amplitude and frequency of the applied pulses. However, tuning these parameters typically induces substantial shifts in the qubit frequencies, thereby imposing significant challenges on the calibration and scalability of large-scale quantum processor architectures. To overcome this limitation, we introduce an approach to control the interaction strength via adjustment of the phase of the parametric modulation. Normally, the parametric phase is utilized in specific quantum simulation protocols for tailoring Hamiltonians [39, 40, 35, 41, 42, 43] or for dynamic on-off switching of couplings under particular resonance conditions [44, 45, 46]. Achieving precise phase control and compensation remains a significant challenge in both quantum gate operations [15, 19, 37] and simulations [35, 38] based on parametric modulation.
@@
In this work, we re-examine the conventional role of the parametric phase, treating it not as an experimental overhead requiring compensation, but as a valuable and versatile quantum control resource. We propose and experimentally demonstrate a new approach where the relative phase between two simultaneous parametric flux pulses—one applied to each of two interacting superconducting transmon qubits—provides a direct control parameter for their coupling strength. A crucial advantage of this phase-based method is that it allows for precise modulation of the interaction strength without inducing the time-averaged qubit frequency shifts that typically necessitate cumbersome recalibration.
@@
We experimentally validate this mechanism in a versatile system of two qubits coupled via a tunable coupler. This configuration allows us to retain the inherent advantages of couplers while simultaneously enhancing the generality and flexibility of our parametric phase-modulation scheme. Our experiments, employing population dynamics and spectroscopy methods, systematically characterize this phase-controlled interaction at both sweet and off-sweet spots, thereby demonstrating its broad applicability under diverse operating conditions.
@@ II.1 Parametric modulation
In practical frequency-tunable transmon qubits, the relationship between the qubit frequency and externally applied flux biases of qubits is nonlinear. An arbitrary waveform generator (AWG) generates a programmed voltage pulse, $V(t)=\bar{V}+\tilde{V}\cos(\omega_{p}t+\phi_{p})$, which is applied to the qubit, and the modulation of the flux bias pulse is described as
@@
where the flux oscillates around the parking flux bias $\bar{\Phi}$ with parametric amplitude $\tilde{\Phi}$, frequency $\omega_{p}$, and phase $\phi_{p}$. The qubit frequency becomes time-dependent and can be expressed as the Fourier series
@@
where the Fourier coefficients are given by
$f_{k}=\frac{\omega_{p}}{\pi(1+\delta_{k,0})}\int_{0}^{{2\pi}/{\omega_{p}}}dt\cos[k(\omega_{p}t+\phi_{p})]\omega(t)$
. The frequency can be approximated as:
@@
when the parking flux bias $\bar{\Phi}$ is set at the sweet spot of the qubit and $\tilde{\Phi}\lesssim\Phi_{0}/2$ (with $\Phi_{0}=h/2e$ being the magnetic flux quantum). Here, $\bar{\omega}$ denotes the time-averaged qubit frequency under modulation, while $\epsilon_{p}$ represents the qubit frequency excursion (i.e., the amplitude of the frequency modulation). If the parking flux bias $\bar{\Phi}$ is set at the off-sweet spot, the frequency can be approximated as
@@
where $\omega_{1,2},~\alpha_{1,2},~b_{1,2}$, $b^{\dagger}_{1,2}$, and $g$ represent the frequencies, anharmonicities, annihilation and creation operators, and bare coupling strength of qubits, respectively. When the parking flux bias $\bar{\Phi}_{1}$ is set at the off-sweet spot with a weak parametric amplitude $\tilde{\Phi}_{1}$, modulating $Q_{1}$ as
$\omega_{1}(t)\approx\bar{\omega}_{1}+\epsilon_{p1}\cos(\omega_{p1}t+\phi_{p1})$
, it induces frequency-modulation sidebands, and the qubit frequency oscillates at many harmonics of the parametric frequency $\omega_{p1}$ (see Appendixes B and C for details). In principle, all parameters $\omega_{1,2},~\alpha_{1,2}$, and $g$ under modulation in the frequency domain become time-dependent due to the interaction [9, 19, 47].
To obtain the effective Hamiltonian in a rotating frame, we define the unitary rotation transformation as
$U=\exp\left(-i\int_{0}^{t}\mathcal{H}_{0}(\tau)d\tau\right).$
For simplicity, we also define $F_{1}(t)=\int_{0}^{t}\omega_{1}(\tau)d\tau$$F_{2}(t)=\int_{0}^{t}\omega_{2}(\tau)d\tau$, $A_{1}(t)=\int_{0}^{t}\alpha_{1}(\tau)d\tau$, and $A_{2}(t)=\int_{0}^{t}\alpha_{2}(\tau)d\tau$. The effective Hamiltonian is then given by
$\begin{split}H_{\mathrm{eff}}=i\frac{dU^{\dagger}}{dt}U+U^{\dagger}(\mathcal{H}_{0}+\mathcal{H}_{I})U=&g(b_{1}b_{2}\exp\{i[-F_{1}-A_{1}(b_{1}^{\dagger}b_{1}-I)-F_{2}-A_{2}(b_{2}^{\dagger}b_{2}-I)]\}\\ &+b_{1}^{\dagger}b_{2}\exp\{i[F_{1}+A_{1}b_{1}^{\dagger}b_{1}-F_{2}-A_{2}(b_{2}^{\dagger}b_{2}-I)]\}\\ &+b_{1}b_{2}^{\dagger}\exp\{i[-F_{1}-A_{1}(b_{1}^{\dagger}b_{1}-I)+F_{2}+A_{2}b_{2}^{\dagger}b_{2}]\}\\ &+b_{1}^{\dagger}b_{2}^{\dagger}\exp\{i(F_{1}+A_{1}b_{1}^{\dagger}b_{1}+F_{2}+A_{2}b_{2}^{\dagger}b_{2})\}),\end{split}$ (6)
where we approximate $g$ as constant. In the $\{|01\rangle,|10\rangle\}$ subspace (indexing the states of coupled qubits $|Q_{1}Q_{2}\rangle$), the above effective Hamiltonian in Eq. (6) is truncated as
$\begin{split}H_{\mathrm{eff}}&=g\exp{i\Delta t}\exp\left({i\frac{\epsilon_{p1}}{\omega_{p1}}[\sin(\phi_{p1})-\sin(\omega_{p1}t+\phi_{p1})]}\right)|10\rangle\langle 01|+\text{H.c.}\\ &=g\sum_{n=-\infty}^{\infty}J_{n}\left(\frac{\epsilon_{p1}}{\omega_{p1}}\right)\exp{i[(\Delta+n\omega_{p1})t+\beta_{n}]}|10\rangle\langle 01|+\text{H.c.},\end{split}$ (7)
where we use the Jacobi-Anger expansion $e^{iz\sin\theta}=\sum_{n=-\infty}^{\infty}J_{n}(z)e^{in\theta}$ (with $J_{n}$ being the $n$-th Bessel function of the first kind), $\Delta=\bar{\omega}_{2}-\bar{\omega}_{1}$, and $\beta_{n}=n(\phi_{p1}+\pi)+\frac{\epsilon_{p1}}{\omega_{p1}}\sin(\phi_{p1})$. The induced effective coupling strength is given by
$g_{\text{eff}}^{n}=gJ_{n}\left(\frac{\epsilon_{p1}}{\omega_{p1}}\right),$ (8)
when the resonance condition $\Delta+n\omega_{p1}=0,n\in\mathbb{Z},$ is satisfied.
@@ II.2 Parametric phase modulation
Here, we observe that the parametric phase $\phi_{p1}$ does not affect the effective coupling strength $g_{\mathrm{eff}}^{n}$ in Eq. (8). However, introducing an additional modulation at $Q_{2}$ with a weak amplitude $\tilde{\Phi}_{2}$, given by
$\omega_{2}(t)\approx\bar{\omega}_{2}+\varepsilon_{p2}\cos(\omega_{p2}t+\phi_{p2}),$
renders the relative phase $\delta\phi_{p}=\phi_{p1}-\phi_{p2}$ as a modulating factor for the effective coupling strength. According to Eq. (6) and (7), the new phase-tunable coupling strength is
@@
We use sum-to-product identities followed by the Jacobi-Anger expansions to simplify the above expression in Eq. (9), rather than applying the separate Jacobi-Anger expansion [39, 40, 35, 46, 42, 43]. The coupling strength in Eq. (9) becomes time-independent
@@
when $\omega_{p}=\omega_{p1}=\omega_{p2}$ and $\Delta+n\omega_{p}=0,n\in\mathbb{Z}$, which results in
$\begin{split}A=&\mathrm{sgn}\left[-\sin\frac{\delta\phi_{p}}{2}\right]\\ &\times\sqrt{\left(\frac{\epsilon_{p1}}{\omega_{p}}-\frac{\epsilon_{p2}}{\omega_{p}}\cos\delta\phi_{p}\right)^{2}+\left(\frac{\epsilon_{p2}}{\omega_{p}}\sin\delta\phi_{p}\right)^{2}}.\end{split}$
Additionally, the coupling strength $g_{\text{phase}}^{n}$ depends not only on the parametric amplitudes $\epsilon_{p1},\epsilon_{p2}$ and frequencies $\omega_{p1},\omega_{p2}$, but also on the phases $\phi_{p1},\phi_{p2}$. This indicates that the parametric relative phase $\delta\phi_{p}$ can be modulated to adjust the argument $A$ of the $n$th Bessel function, thereby modulating the coupling strength. The range of $|A|\in\left[\left|{\epsilon_{p1}}/{\omega_{p1}}-{\epsilon_{p2}}/{\omega_{p2}}\right|,{\epsilon_{p1}}/{\omega_{p1}}+{\epsilon_{p2}}/{\omega_{p2}}\right]$ determines the strength range, which is analogous to interference effects. When the parking flux bias $\bar{\Phi}$ is set at the sweet spot with the parametric amplitude $\tilde{\Phi}\lesssim\Phi_{0}/2$, the form of the effective coupling remains the same as in Eq. (9) but with the parametric frequency $\omega_{p1}~(\omega_{p2})$ and phase $\phi_{p1}~(\phi_{p2})$ doubled.